### v1

# Universal Framework to Evaluate Automotive Perception Sensor Impact on Perception Functions

Abstract
Current research on automotive perception systems predominantly focusses on either improving the sensors for data quality or enhancing the performance of perception functions in isolation. Although automotive perception sensors form a fundamental part of the perception system, value addition in sensor data quality in isolation is questionable. However, the end goal for most perception systems is the accuracy of high-level functions such as trajectory prediction of surrounding vehicles. High-level perception functions are increasingly based on deep learning (DL) models due to their improved performance and generalisability compared to traditional algorithms. Innately, DL models develop a performance bias on the comprehensiveness of the training data. Despite the vital need to evaluate the performance of DL-based perception functions under real-world conditions using onboard sensor inputs, there is a lack of frameworks to facilitate systematic evaluations.
This paper presents a versatile and cost-effective framework to evaluate the impact of perception sensor modalities and parameter settings on DL-based perception functions. Using a simulation environment, the framework facilitates sensor modality testing and parameter tuning under different environmental conditions. Its effectiveness is demonstrated through a case study involving a state-of-the-art surround trajectory prediction model, highlighting performance differences across sensor modalities and recommending optimal parameter settings. The proposed framework offers valuable insights for designing the perception sensor suite, contributing to the development of robust perception systems for autonomous vehicles.
Keywords
Automated Vehicles, Perception Sensors, Sensor Optimisation, Simulation, Trajectory Prediction
I Introduction
Deciding the best approacheciding the optimal perception sensor suite for an assisted and automated driving (AAD) vehicle remains an important, however, highly challenging task for the vehicle manufacturers. This is due to the dependency of most AAD functions on the availability of complete inputs for error-free operation. Recent state-of-the-art (SOTA) algorithms for high-level perception functions such as trajectory prediction and manoeuvre classification are increasingly based on deep learning (DL) models [13, 15, 25]. Most such models are developed and trained using ground truth datasets with no missing data. DL models have an inherent bias on the scope of the training data used during the development. Therefore, the quality and completeness of the inputs to these models play a crucial role in their successful deployment, questioning the performance integrity when such models operate using incomplete input data from onboard perception sensors.
Although the outputs of high-level perception algorithms directly impact the safe operation of the AAD features, research to evaluate the onboard perception sensor configurations that optimise AAD functional performance is almost non-existent [4, 21, 26]. The likely reason is the absence of suitable data to facilitate such studies since similar data logged using different sensor modalities are not readily available in most public automotive datasets [16]. In addition, the parameter settings of the perception sensors remain fixed during the data collection of naturalistic automotive datasets.
The existing approaches to evaluate and optimise the perception sensor suite for an AAD vehicle are threefold [3, 17, 21, 28]
  1. (i)
    Install as many and as high-spec perception sensors using prior engineering experience
  2. (ii)
    Install the sensors as specified by the AAD algorithm developers
  3. (iii)
    Statistical approaches based on existing sensor implementations on autonomous vehicle platforms
Neither approach supports the vehicle designers in selecting an optimal and cost-effective sensor suite for the vehicle platforms since the above approaches are either dependent on tier 1 supplier recommendations or based on empirical knowledge of the design engineers.
To address this knowledge gap, a universal framework developed using a virtual simulation environment is proposed, which enables exploring the impact of a given perception sensor configuration insert on a DL-based perception function. The framework offers maximum flexibility by adopting a modular structure, whereby virtual scenarios can incorporate a multitude of options. These include changes to the environmental conditions relating to both weather changes and road geometries as graphically illustrated in Fig. 1. The perception sensor configuration can be regarded as technology-agnostic, however, for generating synthetic data to conduct case studies, different sensor models can be used.
To demonstrate the utility of the proposed framework, the impact of the radar and the camera sensors is evaluated on a DL-based trajectory prediction model and the optimal setting for a common sensor parameter is determined. Although the case study focuses on surrounding vehicle trajectory prediction, the proposed framework is extendable universally towards other high-level perception functions such as manoeuvre classification and situational awareness.
Figure 1: Proposed Framework
The contributions of this paper are as follows.
  • Presents a universal framework to evaluate the performance of perception functions based on different perception sensor configurations and different environmental conditions.
  • Offers a data-driven approach to specify the optimal perception sensors for an AAD function at the early stages of vehicle development.
  • Demonstrates the use of the proposed framework via a case study, where it evaluates the performance impact of different sensor modalities and parameter settings on a SOTA trajectory prediction model.
In addition, the ‘synthetic data generation’ stage of the proposed framework can produce data to improve the diversity of training and validation datasets. Existing real-world datasets can be augmented with synthetic data to introduce challenging environmental conditions or to compensate for the missing data. Such extended datasets can facilitate retraining models to uncouple domain biases, resulting in robust overall performances.
II Related Works
Vehicles equipped with AAD features tend to have several established perception sensor modalities onboard such as radar, LiDAR, and vision sensors [19, 27]. The number of sensors and sensor modalities on a vehicle is likely to increase in conjunction with the ascending levels of automation defined by the Society of Automotive Engineers (SAE)[14]. Research on optimising an AAD perception sensor suite with or without linking to the performance of the perception algorithms is scarce[17]. It involves considering different perception sensor technologies and sensor parameters such as the range and the field of view (FOV), alongside factors including the environmental conditions applicable to the relevant operational design domain (ODD).
Existing studies on optimising a vehicle’s perception sensor suite mostly concentrate on exploring a single sensor modality as opposed to comparing different modalities [24]. Most studies refer to ‘Perception sensor optimisation’ as the optimal placement of the available sensors to ensure maximum coverage of a predefined area. These approaches also lack the cost-benefit analysis of the impact of the sensor data towards the target perception functions [2, 24].
T. Meng et al. present a framework for automatic selection and arrangement of the perception sensors using a genetic algorithm to solve for the optimal design of the sensor suite. The approach involves several prerequisites including the need to define optional sensors and positioning locations based on experience and feasibility, identify a critical point set (CPS) for emphatic detection, and predefine known ‘constraints’ based on prior engineering experience of the AAD feature. The CPS is based only on known edge cases, limiting the safety optimisation of the solution. Although the prerequisites intend to simplify the complexity of the genetic algorithm and the number of iterations, they act as a limitation demanding prior knowledge. This limits their approach to low SAE level features such as automatic park assist, which are commonly established using imperative algorithms. The opaque nature of DL-based algorithms makes their method unsuitable. [22].
In [11], the authors introduce a simulation-based evaluation methodology to optimise the perception sensor setup in AAD vehicles. They use two virtual driving scenarios to demonstrate the usefulness of the method by analysing the impact of two different radar setups on low-level perception functions: object detection and tracking. The case study is limited to radars and lacks comparison between different sensor modalities. Instead of using sensor models to capture data, they use a development tool for environment perception applications, Baselabs Create, to synthesise the sensor data from the ground truths, i.e. no real perception sensor measurements or virtual sensor models were used during the simulation of the driving scenarios. Also, the sensor parameters were fixed during the synthesis process. Hence, their framework lacks the option to evaluate the sensor suite based on sensor parameters. The evaluation is only based on sensor mounting positions. In a follow-up publication, the authors apply their methodology to explore sensor positioning, specifically by altering the mounting heights to maximise the surround view coverage. However, again the study is limited to radars [12].
In [9], the authors evaluate the perception performance on lane recognition as a function of the sensor data quality. They use different combinations of imaging sensors and camera lenses on lane recognition. However, this study was limited to cameras, and the performance analysis was based on detecting a single static element in the driving environment, ‘lane recognition’.
An established generic technique to optimise the overall sensing capability of a vehicle is sensor fusion. Many sensor-fusion algorithms have been published to leverage the individual strengths of different sensor modalities whilst compensating for each other’s weaknesses to optimise AAD functional performance [6, 1, 23]. However, the objective of most of these algorithms is to offer enhanced coverage and availability of sensor data; therefore, they lack direct associations with the requirements of the deployed perception algorithms, often leading to sub-optimal sensor configurations.
T. Ma et al. introduce a novel metric to evaluate the sensor configurations of automated vehicles using both the perception algorithm performance and the selection of the sensors. Their evaluation metric, ‘perception entropy’, is calculated using Bayesian possibility theory to estimate the uncertainty of the detected targets and is formulated as a Gaussian distribution, conditioned on the sensor configurations. They link the definition of the metric to a selected perception algorithm, hence a change to the algorithm results in a change to the relevant sensor measurements and the relation to the algorithm’s performance.
Another avenue of related research is adaptive sensor weighting techniques based on algorithmic outputs [18, 20]. This is where dynamically varying levels of importance are assigned to different sensor inputs based on the performance of the perception algorithms. Although a promising approach, the technique is applied after making the design decisions for the sensor suite. It also poses challenges in terms of accurately estimating the online algorithmic performance and stabilising the rapid changes in weights assigned to different sensors.
In our view, optimising the sensor suite requires selecting the appropriate sensor modalities and configuring the sensor parameters to realise the optimal performance of the onboard perception algorithms. The sensors currently available in the market offer a range of specifications, often increasing in price with broader operational ranges. Therefore, the ability to perform impact analyses between different sensor configurations and the algorithmic performance of the perception functions is required. The proposed framework is a promising way to achieve the same and reshape the sensor optimisation landscape.
III Problem Formulation
The perception sensor optimisation framework is formulated as a modular structure that brings together the essential elements for investigating the correlations between the configuration of the perception sensors and the performance of a given AAD perception function.
This section details the proposed framework using a simplified illustration of the vehicle’s sensing and perception system, specific to the trajectory prediction of surrounding vehicles as shown in Fig. 2.
Figure 2: Sensing and Perception System - Graphical Representation
The sensing system is a combination of automotive perception sensors and is represented as a function $S$, with internal parameters $r$, and real-world observations, i.e., the ground truths $GT$. The internal sensor parameters represent the sensor specifications such as the range, the resolution, and the FOV.
The ground truths are represented as follows.
$GT=\ \left\{{gt}_{1},{gt}_{2},.....{gt}_{n}\right\}$ (1)
where ${gt}_{1}$ is the track of the target vehicle and ${gt}_{2},.....{gt}_{n}$ are all the other traffic vehicles within a predefined boundary representing the inputs for the prediction model. Similarly, if the sensor-based data is represented as $SD$,
$SD=\ \left\{{sd}_{1},{sd}_{2},.....{sd}_{m}\right\}$ (2)
where ${sd}_{1}$ is the track of the target vehicle and ${sd}_{2}$ to ${sd}_{m}$ are the tracks of detected traffic vehicles within the perimeter by the perception sensor under study, $S$.
Note that not all traffic vehicles may get detected by a given sensor. I.e. $n\geq m$
${gt}_{i}\in\ {GT}$ and  ${sd}_{i}\in\ SD$ where ${gt}_{i}/{sd}_{i}\rightarrow\left(a,b,c,\ldots\right)$.
The parameters $a,\ b,\ c$  can include positional data $(x,\ y,\ z)$, rotational data (yaw, pitch, roll), temporal vectors of the vehicles (velocity, acceleration) etc. depending on the input data requirements of the selected perception function.
If the duration of a vehicle track is from $t=0$ to $t=t$, since the selected trajectory prediction model for this study only requires the positional $x,\ y$ co-ordinates of the vehicles, ${gt}_{i}$ can be represented as follows.
${gt}_{i}=\ \left\{\left(x_{i},y_{i}\right)_{t=0},\left(x_{i},y_{i}\right)_{t=1},.....,\left(x_{i},y_{i}\right)_{t=t}\right\}$ (3)
Similarly,
${sd}_{i}=\ \left\{\left(x_{i},y_{i}\right)_{t=0},\left(x_{i},y_{i}\right)_{t=1},.....,\left(x_{i},y_{i}\right)_{t=t}\right\}$ (4)
As shown in Fig. 2, the sensor captures the ground truth data of a traffic object and outputs the sensor measurement, $M$, based on the sensor parameter $r$.
$M=S\left(GT,r\right)$ (5)
The detection function, $D$ outputs the detected traffic, $SD$, using the sensor measurements $M$ as the input, where $b$ represents internal parameters of the detection function. Therefore,
$SD\ =\ D\left(M,\ b\right)$ (6)
By substitution from (5),
$SD\ =\ D\left(S\left(GT,r\right),\ b\right)$ (7)
Since the detection algorithm deployed in an automated vehicle is fixed, the parameter values, $b$, are also fixed for the application. Hence (7) can be simplified to:
$SD\ =\ \hat{D}\left(\hat{S}\left(GT,r\right)\right)$ (8)
Forming a composite function, $F$:
$SD\ =\ F\left(GT,\ r\right)$ (9)
Equation (9) exemplifies the direct correlation between object detection and the technical parameters of a perception sensor. Therefore, to evaluate how a sensor parameter affects the output of the trajectory prediction model, one can compare the performance (output accuracy) of the model based on the sensor data with different sensor parameter values $(r)$ vs the performance of the model based on the ground truths.
As pictorially illustrated in Fig. 2, the trajectory prediction model outputs a probability distribution $P\left(Y\middle|X\right)$ over the future co-ordinates of the target vehicle’s predicted trajectory represented as $Y_{pred}$.
To derive a set of predictions eliminating the impact of the sensor parameters, one can use the ground truths as the input data and use the results as the baseline to compare the impact a sensor parameter has on the model’s output. If the predicted trajectory vector based on ground truths is $Y_{pred}^{GT}$, and the predicted trajectory vector based on a preset value of $r$ for the ’sensor parameter under-study’ is $Y_{pred}^{r}$, where,
$Y_{pred}^{GT}=\ \left[y_{t=1}^{GT},\ldots,y_{t=f}^{GT}\ \ \right]$ (10)
and
$Y_{pred}^{r}=\ \left[y_{t=1}^{r},\ldots,y_{t=f}^{r}\ \ \right]$ (11)
for a prediction time horizon of $f$ time-frames to the future, the error between the two trajectory predictions based on the ground truths $(GT)$ and the sensor detected traffic objects $(SD)$, can be shown as,
${E\ }_{RMSE}^{r}\ =\ \sqrt{\frac{1}{f}\ \sum_{i=1}^{f}\left(\ y_{t=i}^{GT}-\ y_{t=i}^{r}\right)^{2}}$ (12)
Where ${E\ }_{RMSE}^{r}$ is directly correlated to the sensor parameter setting, $r$.
Thereby determining the optimal value for each sensor parameter, $r$, which minimises the error, ${E\ }_{RMSE}^{r}$, one can optimise the configuration of the perception sensor suite. With a specified range of values being from $i$ to $j$ for given perception sensor parameter,
${arg\ \ {\min_{r\ \in\ \left(i,j\right)}}}({{E\ }_{RMSE}^{r})}$ (13)
The interval $\left(i,j\right)$ depends on the manufacturer’s specifications of a given sensor for a selected ODD of an AAD function.
IV Methodology – Proposed Simulation Framework
Fig. 1 illustrates the proposed framework which implements the modular structure formulated in section III consisting of four stages: (i) synthetic data generation, (ii) data pre-processing, (iii) data utilisation, and (iv) data analysis.
IV-A Synthetic Data Generation
A virtual environment is used to simulate the required driving scenarios for data generation. Following an empirical study, IPG CarMaker (CM) is chosen as the simulation tool for the framework due to its widespread use in the industry [10].
Figure 3: Synthetic Data Generation Setup
The proposed framework offers maximum flexibility in choosing over a multitude of variable settings under two avenues. As shown in Fig. 3, firstly, the user can select the type of perception sensor and set the sensor parameter configuration in line with a test strategy. Secondly, they can select different vehicle types, manoeuvres, and environmental conditions when creating the scenarios, allowing them to incorporate edge cases, as necessary.
Execution of the tests and data logging is automated using a CM test automation feature, Test Manager. The simulation data is saved in a results file in the ERG file format.
IV-B Data Preprocessing
Customised data preprocessing scripts are created in MATLAB for each sensor modality due to the differences in the recordable quantities for different sensor modalities. The script identifies the surrounding vehicles within a predefined boundary for the target vehicle and derives the input features for each ’vehicle of interest’ for the trajectory prediction model. It encapsulates the instantaneous road angle at each time frame to enable converting the data between the ego-centric frame of reference and the target vehicle’s frame of reference and generates the input data files for the trajectory prediction model. The ground truth trajectory for the target car is obtained directly from the simulation tool.
The data pre-processing MATLAB scripts are available in GitHub at https://github.com/asha-gamage/perception4prediction
IV-C Data Utilisation
The ‘Utilise Data’ stage represents the use of the perceived data from the driving environment as the inputs to the prediction model to obtain inferences. Since a trajectory prediction model is used, two probability distributions of the predicted trajectory of the target vehicle are obtained, one based on the ground truths and the other based on the sensor data.
IV-D Data Analysis
The metric root mean squared error (RMSE) is used for analysing the data. The RMSE values calculated on the predicted trajectories from the model using both the ground truth data and the sensor data are plotted against the prediction time horizon to compare the impact of different sensor modalities and parameter settings.
V Case Study: Surround vehicle Trajectory Prediction
The merits of the framework are demonstrated using a DL-based model for trajectory prediction of surrounding vehicles. Six driving scenarios are created with a target vehicle performing a lane change manoeuvre on a highway. Lane change is identified as one of the safety-critical manoeuvres causing a high number of road accidents on highways [7, 8]. Therefore, developing knowledge on the impact of perception sensor configuration on lane change prediction performance is of high value.
V-A Model Selection
After comparing several SOTA DL-based vehicle trajectory prediction models for their suitability for highway driving, a model proposed by N. Deo et al. is chosen [4]. This model predicts future vehicle motion by incorporating manoeuvre-based predictions and utilising convolutional social pooling. The manoeuvre-based approach categorises vehicle trajectories into distinct manoeuvres, whereas the convolutional social pooling helps capture spatial dependencies between the surrounding vehicles’ trajectories more effectively. The model, referred to as the CS-LSTM, combines the benefits of both the convolutional neural networks (CNN) and long short-term memory networks (LSTM) architectures and has been trained using a large naturalistic dataset, NGSIM, sourced from the US highways [5].
The architecture of the chosen model consists of an LSTM encoder and convolutional social pooling layers followed by an LSTM decoder.
V-B Inputs and outputs
Inputs to the model are the track histories of the target vehicle and its surrounding vehicles, taken for the time frames from $t=t-t_{h}$ to $t=t$ and consist of the $x,\ y$ co-ordinates as detailed in [4]. For our analysis, the track histories are derived in two ways.
a) Sensor-based data: The longitudinal and lateral displacements of the detected traffic objects in the sensor frame of reference are used to calculate the track histories in the form of Cartesian co-ordinates.
$X^{sen}=\ \left[x_{t-t_{h}}^{sen},\ldots,x_{t}^{sen}\ \ \right]$ (14)
where,
$x_{t}^{sen}=\ \left[{\hat{x}}_{t}^{tar},\ {\hat{y}}_{t}^{tar},\ {\widehat{\ x}}_{t}^{sur_{1}},\ {\widehat{\ y}}_{t}^{sur_{1}},\ldots,{\widehat{\ x}}_{t}^{sur_{n}},\ {\widehat{\ y}}_{t}^{sur_{n}}\ \right]$
are the $x,\ y$ co-ordinates of the target vehicle $(tar)$ and $n$ number of surrounding vehicles $(sur_{n})$ at time frame, $t$.
b) Ground truth data: The ground truth Cartesian co-ordinates for the traffic objects are directly read from the simulation platform as the global positions of each traffic object.
$X^{GT}=\ \left[x_{t-t_{h}}^{GT},\ldots,x_{t}^{GT}\ \ \right]$ (16)
where,
$x_{t}^{GT}=\ \left[{\hat{x}}_{t}^{tar},\ {\hat{y}}_{t}^{tar},\ {\widehat{\ x}}_{t}^{sur_{1}},\ {\widehat{\ y}}_{t}^{sur_{1}},\ldots,{\widehat{\ x}}_{t}^{sur_{n}},\ {\widehat{\ y}}_{t}^{sur_{n}}\ \right]$
Positional co-ordinates from both approaches are converted to the target vehicle’s frame of reference before using as inputs to the model. The output of the model infers the conditional probability distribution $P\left(Y\middle|X\right)$. Therefore, the inferred future $x,\ y$ positions at each future predicted time-frame are the parameters of a bivariate Gaussian probability distribution, where $\mu^{t}$ is the mean vector and $\Sigma^{t}$ is the covariance matrix.
$\mu^{t}=\begin{pmatrix}\mu^{t}_{x}\\ \mu^{t}_{y}\end{pmatrix},\Sigma^{t}=\begin{pmatrix}({\sigma^{t}_{x}})^{2}&\sigma^{t}_{x}\sigma^{t}_{y}\rho^{t}\\ \\ \sigma^{t}_{x}\sigma^{t}_{y}\rho^{t}&({\sigma^{t}_{y}})^{2}\end{pmatrix}$ (18)
The mean values are used as the inferred $x,\ y$ positional co-ordinates.
V-C Scenario Generation
Three scenario categories are defined with two representing a road segment from the US Highway 101 and the remainder, a straight highway road segment. The scenarios cover the most influential positions for a lane change by a target vehicle to impact the trajectory of the ego car, i.e. a lane change into the same lane as the ego car, directly in front of it, and in front of the lead vehicle to the ego car.
By considering each scenario category for both directions, six driving scenarios are created with a target vehicle performing a lane change manoeuvre. The ego car is represented by a typical passenger vehicle type along with several different vehicle types representing the surrounding vehicles. The ego car is equipped with a radar sensor and a stereo camera sensor.
  • Scenario 1: Lane change to the Right - In front of the ego car (US101)
  • Scenario 2: Lane change to the Left - In front of the ego car (US101)
  • Scenario 3: Lane change to the Right - In front of the lead car (US101)
  • Scenario 4: Lane change to the Left - In front of the lead car (US101)
  • Scenario 5: Lane change to the Right – In front of the lead car (Straight road)
  • Scenario 6: Lane change to the Left - In front of the lead car (Straight road)
The highway road segment consists of three lanes in each driving direction. Supplementary Data Tables A.1 and A.2 illustrate scenarios 1 and 2 with detailed descriptions in line with the definition of a scenario in [29].
Figure 4: Mounting Positions of Sensors
Source: IPG CarMaker
High-fidelity sensor models for the radar and the stereo camera from CM are mounted on the ego vehicle positioned as shown in Fig. 4. The radar uses signal-to-noise ratio for detecting the surrounding objects with consideration given to the propagation and atmospheric losses whereas the camera uses a bounding box of the object for calculating the occlusions to determine whether an object is detected. The detection confidence is compared with a predefined threshold, which for this study is set to 0.6. Tables I and II list the default sensor specifications for the radar and the camera sensors except for the technical parameter being changed in the study, horizontal FOV (HFOV) and the range of the sensor, which is fixed at 100 m.
TABLE I: Radar Sensor Parameters
Technical Parameter Value
Accuracy Resolution
Distance (m) 0.4 1.8
Azimuth (deg) 0.1 1.6
Speed (km/h) 0.1 0.4
Cycle time (ms) 60
Transmit frequency (GHz) 77
Transmit Power (dBm) 14
Noise bandwidth (Hz) 25000
Noise figure (dB) 4.8
TABLE II: Camera Sensor Parameters
Technical Parameter Value
Horizontal Vertical
Resolution (px) 1280 960
Resolution factor [-] 50
Camera type Stereo
Baseline (m) 0.12
Disparity error (px) 0.5
Focal length (px) 2400
VI Results and Analysis
A test run is established using a predefined scenario with one sensor modality and the sensor’s HFOV set as a variable. It is then simulated iteratively varying the assigned value to the variable, HFOV, as listed in Table III. Identical sets of simulation test runs are conducted for the radar and the camera sensors separately.
TABLE III: Test Run Definitions - HFOV Variation
Relative Velocity Scenario LC01: Lane Change to the Right Scenario LC02: Lane Change to the Left
Range = 100 m VFOV = ${15}^{\circ}$ Range = 100 m VFOV = ${15}^{\circ}$
HFoV Test ID HFoV Test ID
$V_{R}$ = 5 km/h ${30}^{\circ}$ LC01_RelV5_30 ${30}^{\circ}$ LC02_RelV5_30
${60}^{\circ}$ LC01_RelV5_60 ${60}^{\circ}$ LC02_RelV5_60
${90}^{\circ}$ LC01_RelV5_90 ${90}^{\circ}$ LC02_RelV5_90
${120}^{\circ}$ LC01_RelV5_120 ${120}^{\circ}$ LC02_RelV5_120
${150}^{\circ}$ LC01_RelV5_150 ${150}^{\circ}$ LC02_RelV5_150
${180}^{\circ}$ LC01_RelV5_180 ${180}^{\circ}$ LC02_RelV5_180
VI-1 Impact of the Radar sensor’s ‘HFOV’ on Trajectory Prediction
Figs. 5, 6 and 7 show the derived RMSE values based on the radar sensor plotted against the prediction horizon of 5 s for the three scenario categories.
Figure 5: Impact on RMSE - Changing Radar’s HFOV: RLC Vs LLC - In front of Ego (US101)
Figure 6: Impact on RMSE - Changing Radar’s HFOV: RLC Vs LLC - In front of Lead (US101)
Figure 7: Impact on RMSE - Changing Radar’s HFOV: RLC Vs LLC - In front of Lead (Straight Road)
It is observed that the lowest setting of ${30}^{\circ}$ for the HFOV has produced the RMSEs closest to the ground truth-based RMSEs in all three scenario categories explored: (i) in front of the ego car on US101, (ii) in front of the lead car on US101 and (iii) in front of the lead car on the straight road. The intuitive expectation is for the sensor-based data obtained with a larger HFOV to perform better when used as the input to the prediction model. However, the results exemplify that the RMSEs of the predicted trajectories are optimal when the measurements are taken with the lowest HFOV of ${30}^{\circ}$.
It is believed since the dominant manoeuvre of all six scenarios is a lane change into the ego car’s lane, the narrow focus given by a smaller HFOV of the radar sensor enables the DL model to make a better prediction by reducing the complexity and unnecessary noise. Since for a lane change, the highest impact on the target vehicle is imposed by the vehicles in its closest proximity and directly ahead, a narrow HFOV improves focus by reducing the irrelevant data of unrelated surrounding traffic. Another likely rationale is that the data captured at ${30}^{\circ}$ HFOV is a close representative of the relevant area of interest around the target vehicle for the prediction model. The CS-LSTM model uses the BEV representation of the scene with a predefined boundary around the target car. Therefore, when the sensor data is based on a wider HFOV, the inputs may appear distorted reducing the model’s performance.
The above observations highlight the significance of determining the value addition by deploying high-specification perception sensors within the sensor suite of an AAD vehicle as mentioned in [17]. Although the expectation is that high-specification sensors would lead to high-quality data, the results of this experiment indicate that the performance of DL models might not always positively correlate with data from high-specification sensors.
VI-2 Impact of the Camera sensor’s ‘HFOV’ on Trajectory Prediction
The derived RMSE values based on the camera sensor plotted against the prediction horizon for the three scenario categories are shown in Figs. 8, 9 and 10.
Figure 8: Impact on RMSE - Changing Camera’s HFOV: RLC Vs LLC - In front of Ego (US101)
Figure 9: Impact on RMSE - Changing Camera’s HFOV: RLC Vs LLC - In front of Lead (US101)
Figure 10: Impact on RMSE - Changing Camera’s HFOV: RLC Vs LLC - In front of Lead (Straight Road)
In comparison to the observations made with the radar-based data, the gap between the RMSEs from the camera-based data and the ground truth RMSEs is much larger. It suggests that in general radar is likely to be a better-suited sensor modality for surrounding vehicle trajectory prediction in comparison to the camera. However, it is acknowledged that much more extensive data is required to solidify such a presumption.
The camera sensor is known to struggle capturing rapid or complex movements accurately and this is likely the case for the target car and its surrounding vehicles during lane changes, especially with the cut-in manoeuvres included in the scenarios. The camera is also highly susceptible to occlusions. Figs. B.1, B.2, and B.3 in Supplementary Data B illustrate the detection of the target and the relevant surrounding vehicles between the two sensor modalities. Fig. B.2 shows how the occlusions erroneously elongate the duration of the lane change with the camera sensor, whereas, Fig. B.3 shows for the straight road scenarios, the algorithm even fails to identify the end of the lane change manoeuvres due to the camera failing to detect the target car following the lane change due to being occluded by the lead vehicle.
For the lane changes to the right in scenarios ‘In front of the ego’ and ‘In front of the lead’, a HFOV setting of ${120}^{\circ}$ seems to provide the closest RMSEs to the ground truth RMSEs. However, for the left lane changes (LLCs) in those scenarios, the HFOV settings do not show a recognisable trend in terms of producing the RMSEs closest to the ground truth equivalents. The gap between the two RMSEs also remains considerably high towards the latter part of the prediction horizon.
With the ‘StraightRd’ scenarios, for the right lane changes (RLCs), a HFOV of ${180}^{\circ}$ seem to produce the RMSEs closest to the base RMSEs whereas, for the LLCs, the closest to the base RMSEs is produced with a HFOV of ${120}^{\circ}$.
However, with the two scenarios based on the US101 road segment, for the RLCs, a ${120}^{\circ}$ HFOV seems to work the best capturing all the relevant data without losing considerable resolution. However, the LLCs are inclined to be complex since they are manoeuvres into a high-speed lane and lack identifiable correlations with the HFOV changes. The considerably high RMSE gap is likely to be due to the camera’s susceptibility to dynamic scene changes and occlusions. The discrepancy in the optimal HFOV settings between the straight road and the US101-based scenarios is likely due to the differences in the road geometry.
The common observation between the RLCs and the LLCs for all three scenario categories with the camera sensor data is that the HFOV of ${30}^{\circ}$ produces the largest gap between the two RMSE curves and the variation of the RMSE curves amongst the other HFOV settings not having such a considerable impact. I.e. for the camera sensor, varying FOV does not appear to impact the trajectory prediction model performance to a noticeable extent. The likely explanation for the above is the camera’s inability to capture critical inputs at the lowest HFOV of ${30}^{\circ}$, i.e. the largest gap between the RMSEs based on ground truths and the camera data is with a ${30}^{\circ}$ HFOV setting.
Camera seems to require a broader view to compensate for the lack of visual data, especially with the lateral movements of traffic that happen during a lane change manoeuvre. On the other hand, lack of impact when increasing the camera’s HFOV beyond ${60}^{\circ}$ alludes to its ability to detect and process the relevant traffic objects quite robustly beyond a certain minimal HFOV such that further increases result in redundant data.
Overall, the noticed difference in performance of the trajectory prediction model when the input data is sourced from two different sensor modalities is likely due to the fundamental operational differences between the two modalities. It is believed that the divergent strengths of the two sensor modalities, such as the radar’s ability to focus on the relevant traffic objects with a narrow HFOV whereas the camera’s dependency on visual data demanding a broader view to capture the required data, causes these observations. In addition, the high sensitivity of the camera to occlusions compared to the radar’s robustness to such environmental factors is a likely factor contributing to the difference. Hence, for an AAD vehicle deploying a trajectory prediction model such as the CS-LSTM, the inherent characteristics of the radar yield itself as an essential constituent of the perception sensor suite.
Another interesting observation is that with the radar sensor inputs the prediction model seems to perform better for the RLC manoeuvres in comparison to the LLC manoeuvres. By comparing the radar’s detections of the target vehicle and its surrounding vehicles between the bi-directional lane changes as shown in Fig. C.1 in Supplementary Data C, it is seen that the target car and the surround car, NIO are detected less consistently in the LLC scenarios compared to the RLC scenarios. This explains the observation of slightly better prediction performances for the RLC manoeuvres since longer segments of missing data with the LLC manoeuvres contribute to poor predictions.
As a crude litmus test to check the influence of missing data on the prediction model’s performance, data imputation is applied to the original radar data using linear interpolation. It is noted that with compensated data, regardless of the direction of lane change, the corresponding RMSEs along the prediction horizon are compatible with each other as shown in Supplementary Data D Figs. D.1 and D.2.
In all scenarios studied, the RMSEs improve following data imputation, which makes for the case that the performance of the CS-LSTM degrades considerably with impaired detection of the target car and its surrounding vehicles. This is understandable since the loss of detection, even for a short duration, directly impacts the determination of the track history, which is the main input used by the DL model for making predictions. This highlights the need for further research to explore the robustness of the SOTA DL trajectory prediction models when deployed to operate using incomplete real-world data.
VII Conclusion
Insights on how different types of perception sensors and their parameter configurations contribute to optimising the perception sensor suite of an AAD vehicle are highly desirable as the automotive industry progresses towards full automation. However, there are hardly any studies to understand the correlations between different perception sensor modalities/ sensor parameter settings and the performance impact on a DL-based perception function such as trajectory prediction. Such knowledge is essential to optimise the overall perception system efficiently and cost-effectively. It is identified that the lack of studies is mainly due to the challenge of obtaining sensor data with varying parameter settings whilst the rest of the variables remain constant in a highly dynamic driving environment.
As a way forward, a framework is introduced which combines a virtual simulation environment with sensor models to generate synthetic data for different sensor modalities. The framework involves all stages of the perception pipeline and a data analysis stage, offering maximum flexibility with its modular approach. Therefore, it can be used to conduct a multitude of different sensitivity studies. It facilitates quantitative sensitivity analyses between sensor configurations and the performance of a perception function based on an applicable metric to the selected perception function. Overall, the framework facilitates optimising the perception sensor suite by tailoring the sensor configuration to the needs of the perception algorithms implemented on the vehicle.
The contribution of our framework is demonstrated by using it to investigate the optimum parameter setting for the HFOV, in addition to selecting between the sensor modalities radar and camera for the application. The results suggest that a radar sensor with a narrow HFOV is the most suited sensor configuration for the perception algorithm tested.
The possible uses of the proposed framework to the vehicle manufacturers and the regulatory bodies include,
  1. (i)
    Making efficient design choices of the sensor suite at the early stages of the development cycle
  2. (ii)
    Rapid prototyping of concept vehicle platforms to make informed choices between the sensor suite and the perception algorithms
  3. (iii)
    Serve as a means for cost-benefit analysis on sensor technologies and sensor specifications
  4. (iv)
    Defining minimum safety requirements to be complied with to achieve system homologation
Future work will extend the variety of the scenarios and the sensor parameters being studied. It is also planned to explore the optimal perception sensor selections and parameter configurations for different environmental conditions such as adverse weather and different road geometries.